Operations Report – February 22nd

Crew 263 Operations Report 22-02-2022

SOL: 1

Name of person filing report: Mathéo Fouchet

Non-nominal systems: NA

Notes on non-nominal systems: NA


Spirit rover used: No

Hours: 157.2

Beginning charge: 100%

Ending charge: 100%

Currently charging: handled by Atila

Opportunity rover used: Yes

Hours: 84.4

Beginning charge: 100%

Ending charge: 98%

Currently charging: left uncharged, handled by Atila

Curiosity rover used: Yes

Hours: 168.4

Beginning charge: 100%

Ending charge: 96%

Currently charging: left uncharged, handled by Atila

Perseverance rover used: No

Hours: 228.1

Beginning charge: 100%

Ending charge: 100%

Currently charging: handled by Atila

General notes and comments: NA

Summary of Hab operations:

WATER USE: 13 gallons

Water (static tank): 475 gallons

Water (loft tank): 44 gallons

Water Meter: 01560605 units

Static to Loft Pump used: Yes

Static tank pipe heater: On

Static tank heater On

Toilet tank emptied: No

Summary of internet:

Summary of suits and radios: I have changed the battery from suit #8 to make the suit works. For the other suits, I had a problem with several chargers that didn’t make any light, either green nor red, when plugged. However, by switching the plugs for some of them, the chargers light works again. As I see a green light on the chargers, I assume that the suits are actually charging.

Moreover, the fans from suit #10 are not working. I have just checked the battery voltage which is 13V, I have also checked that the fans can physically move and nothing is blocking them to do so. I would be grateful for any help.

Summary of GreenHab operations:

WATER USE: 8 gallons at 8:30am and 6 gallons at 4pm

Heater: On

Supplemental light: On

Harvest: 1 pepper (6g) and sweetie tomatoes (114g)

Summary of ScienceDome operations:

Dual split: Off

SOC: 100%

Summary of RAM operations: NA

Summary of any observatory issues: NA

Summary of health and safety issues: See HSO report

Questions, concerns and requests to Mission Support: We put a trash in the rear airlock for pickup. Thanks a lot!

Journalist Report – February 22nd

Sol 1 – Searching for a routine, or almost…

Author: Nicolas Wattelle

Yesterday was the day of the firsts, now we need to move on, to evolve, to become real analog martionauts. Although we have not discovered everything yet, things are becoming familiar to us.

We all woke up at 6:45am, in order to be ready for the sport session at 7 am. Our first one together! Valentine had prepared a cardio training, which felt a bit rough to all of us, but we will say that it is normal for a first session (Crew 240 reported similar feelings in Pierre’s -the journalist- report a month ago…).

Anyway, this was a good cohesion moment, a great way to begin the day!

Then, the breakfast was in some sort similar to the previous day, carefully prepared by two members of the crew. It was a hearty meal for the four-brand new martionauts who had to prepare for the challenge of the day: the first crew EVA.

This one had multiples objectives: first, to get used to the suits, the communication protocol, the crew member buddy system… It may seem like it is obviously easy but, be aware that this is not the case, even for people who mentally prepared for a year. Second, to check the status of the atmospheric experiment left on the field by Crew 240: the weather station, the LOAC (an amazing aerosol meter), the electric field mill… There, we had some surprises, and we could not do all of what was planned for this phase of today’s EVA. However, we will put things back in order in the coming days. Last, to do a first on-field trial of the Hololens. I was chosen to wear the glasses (during all the outing) for Augmented Reality (AR) testing. But it did not go well. The glasses fell slowly during the EVA, so much that when this phase began, I was almost half blind… I had to let them completely fall around my neck to get my vision back and we shortened the EVA. Note for the future tests: prepare a better mounting to keep the glasses on. A case to follow!

Lunches are always discussion moments. We speak about a lot of things, such as our “old” terrestrial life or our new Martian life. But lately, the food rationing topic came back frequently: we have three weeks’ worth of food already in the station but how do we manage to correctly feed everyone without being too drastic or too neglecting? On top of that, how to make it taste good every day? Some of us are more likely to be cook than others (guess what, that would not be me…), so they lead the discussion about this puzzle (and I thank them a lot!).

The afternoon was quite similar to the previous one. Everyone was focused on personal experiments in different rooms of the station: ScienceDome, GreenHab, upper and lower deck of the Hab… And it went well!

As the crew journalist, I would like you, dear reader, to come to the backstage of the mission with me, and discover more about my mates I am living this journey with. To give you this glimpse, some days, I may ask some of them what are their thoughts about different subjects. My question of the day is:

Why are you here?

Mathéo did not hesitate for a second, he clearly wants to make a first step into astronauts’ life. He took this MDRS mission as an opportunity to get closer to his dream.

To Marine, research is part of her life, and she aims to make a living out of it. This project is a concrete and original way to get a first step into research, so she has get involved in it!

Léa also has a particularly clear idea of her motivations. The operational aspect of crewed flights is what she is keen on, and analogs are an efficient way to understand the full scope of it!

Elena wants to work in human spaceflight, and in order to be a good an engineer and designer, one must think as an astronaut, and she is taking this mission as an opportunity to walk in astronauts shoes!

Astronomy Report – February 22nd

Name: Marine Prunier

Crew: 263
Date: 22 Feb 2022


Robotic Telescope Requested (choose one): MLC-RCOS16

Objects to be Imaged this Evening: TOI 1555.01 (TESS catalog exoplanet)

Images submitted with this report: None

Problems Encountered: The MLC-RCOS’s dome was probably closed yesterday because the photos of TOI 961.01 are all completely black. No photometry can be performed.


Solar Features Observed: /

Images submitted with this report: /

Problems Encountered: /

EVA Report – February 22nd

Crew 263 EVA Report 22-02-2022

EVA # 1

Author: Cerise Cuny

Purpose of EVA: Initiation EVA, batteries replacement of atmospheric experiments, mapping of station with Augmented Reality, location search for future set up of antenna.

Start time: 9:39

End time: 10:39

Narrative: We checked the rover batteries (all full), and left with Opportunity and Curiosity towards Marble Ritual. We arrived at 9:43, we walked a bit, to see how comfortable we were with the EVA suits, and to get used to them. We took time to admire the Martian landscape. We left Marble Ritual with the rovers at 10:00 to join the atmospheric experiment site. We removed the LOAC because it does not bear snow storms well. We plugged a battery to PurpleAir and re-oriented the anemometer to the North. We had forgotten to bring a garbage bag to protect the battery from the rain ; as it is absolutely necessary to protect it before tonight’s snow storm, we asked our HabCom to put a garbage bag in the air lock. Elena (my buddy) and I went to fetch it while the other two stayed near the weather station to fix some things. We did not lose sight of the other pair and we stayed within the radio range.

When we were done with the atmospheric experiment we went back to the station with the rovers (10:29) and left them there to walk around the station. Nicolas was not able to test the augmented reality to map the MDRS because the HoloLens fell around his neck during the first part of the EVA. We found a nice spot for the communication antenna: far enough from the Science Dome so that it can’t fall on it even in case of extreme winds, but close enough for us to plug its 20m electric cable in the Dome. At 10:39 we were back in the air lock to end the EVA.

Destination: Marble Ritual, atmospheric experience site (at intersection between Cow Dung Road and Entrance Road), and around the station.

Coordinates (use UTM NAD27 CONUS): 518600, 4250700

Participants: Cerise Cuny, Marine Prunier, Nicolas Wattelle, Elena Lopez-Contreras.

Road(s) and routes per MDRS Map: Entrance Road, Cow Dung Road.

Mode of travel: Foot, Rover

Sol Summary – February 22nd

Crew 263 Sol Summary Report 22-02-2022

Sol: 1

Summary Title: First steps on Mars!

Author’s name: Cerise Cuny

Mission Status: Nominal

Sol Activity Summary: At 7:00 AM, we did our first sport session organised by our HSO. After breakfast we prepared for our first EVA ; it was a bit longer than expected because we struggled with putting on the HoloLens inside the Spacesuit’s helmet. During the EVA, my crew mates really felt like they were on Mars and were positively surprised by how the spacesuits felt very real (or at least how they imagine a real Mars experience), because of their weight, the impact of the suit and gloves on mobility, and the constraint of the helmet on vision. We were unable to conduct one of the AR experiments because the HoloLens fell around the neck of Nicolas. Elena will improve the device for future EVAs.

After lunch some of us went through one of Crew 240’s human factors experiments while others did some prep work for tomorrow’s experiments such as the ultrasound protocol and the setup of the antenna. Our Journalist recorded the beginning of his podcast’s first episode, which sounds absolutely amazing! Bread is in the oven for dinner, it smells wonderful in the hab!

Look Ahead Plan: If the terrain conditions allow it, I will go on a second Initiation EVA to Marble Ritual with the three remaining crew members. In the afternoon, all crew members will perform ultrasound sessions, both as operators and subjects, in order to test a CNES protocole.

Anomalies in work: None to report.

Weather: Cloudy.

Crew Physical Status: Perfectly fine.

EVA: One performed, everything went fine except for the aforementioned HoloLens issue.

Reports to be filed: Sol Summary, EVA Request, EVA Report, Operations Report, Journalist Report, Astronomy Report, HSO Report.

Support Requested: Only one GPS is working (we already tried changing batteries for the other ones) ; if possible we would like to have another working one for redundancy.

Copyright © The Mars Society. All rights reserved. | Main Site